Tempura, Inc., is considering two projects. Project A requires an investment of $44,000. Estimated annual receipts for 20 years are $23,000; estimated annual costs are $12,500. An alternative project, B, requires an investment of $77,000, has annual receipts for 20 years of $33,000, and has annual costs of $18,000. Assume both projects have a zero salvage value and that MARR is 13.0 %/year.
Answer #1

Present worth of A = -44000 + (23000 - 12500)*(P/A, 13%,20)

= -44000 + 10500 * 7.024752

= 29759.90

Present worth of B = -77000 + (33000 - 18000)*(P/A, 13%,20)

= -77000 + 15000 * 7.024752

= 28371.27

As Present worth of A is more it should be selected